# Learning Repetition-Invariant Representations for Polymer Informatics

### v1

@@ 4.3 Ablation Study
In this experiment, we evaluate the impact of different training pairs on MeltingTemp (homopolymer) and IP (copolymer). Figure 3 presents the results. Training pair (with 3-RU augmentation) reaches the best improvement. Further increasing the merge size beyond 3 leads to convergence, with no significant additional gains. These observations are highly compatible with our theory in Section 3.3.2: the training set lacks binary-update supervision, and the hyperdegree-2 gradient strength plateaus for merge size .
@@
We investigate the effect of different merge ratios in repeat-unit augmentation on the training set. Since we have demonstrated both theoretically and empirically that a repeat size of three is optimal, this experiment uses the training set . Unsurprisingly, we can observe that in Table 5, a balanced 1:1 ratio gives the best performance, while reducing the proportion of augmented samples to 1:0.9 and 1:0.8 leads to a monotonic decline in accuracy. This observation further validates our analysis: weakening binary-update supervision (from hyperdegree-2 supernodes) slightly diminishes training effectiveness. More ablation studies are provided in Section C.2.
@@ 4.4 Visualization
Figure 2 visualizes the t-SNE embeddings generated by GCN and GRIN, trained with and without augmentation, evaluated on test sets augmented up to 20 RUs. As discussed in Section 1, for the same polymer, GCN’s embeddings remain sensitive to repeat size, even with augmentation. In contrast, as Figure 2 (b), GRIN produces stable embeddings across all test lengths, when combined with 3-RU augmentation collapses all repeats into a single cluster—demonstrating true repetition invariance.
@@ 5 Conclusion
In this work, we addressed limitations of prior polymer learning methods that modeled polymers as single monomers and thus failed to capture their realistic, repetitive structures. We made the first attempt to achieve repetition-invariant representation learning. To this end, we proposed GRIN, which combined MST-aligned max aggregation with a repeat-unit-based data augmentation strategy. We provided a theoretical analysis of the minimal augmentation size required to achieve invariance. Experiments on both polymer and copolymer datasets showed that GRIN learned embeddings that reliably generalized to polymers with long repeat lengths.
@@ Appendix A Dataset Details
We evaluate our models on four homopolymer property datasets and two copolymer datasets, each split into 60% training, 10% validation, and 30% test sets. To prevent data leakage, we apply repeat-unit augmentation independently within each split: for every dataset, we generate augmented graphs with repeat sizes from 2 to 60. On the training set, we benchmark four augmentation schemes—{1,2}, {1,3}, {1,2,3}, and {1,4}—while the test set is evaluated on graphs spanning all repeat sizes from 1 through 60.
@@
GlassTemp, MeltingTemp, PolyDensity, and O2Perm. Targets are predicting glass transition temperature (GlassTemp, ∘C), polymer density (PolyDensity, g/cm3), melting temperature (MeltingTemp, ∘C) and oxygen permeability (O2Perm, Barrer), respectively. The first three datasets are extracted from Polymer Info [22], O2Perm is compiled from the Membrane Society of Australasia portal following Thornton et al. [27].
The distributions displayed in Figure 4 (a)-(d) indicate that GlassTemp (7000 samples) and MeltingTemp (3600 samples) exhibit only moderate skew, which corresponds to stable performance across all models. PolyDensity (1700 samples) is nearly Gaussian but limited to its smaller size, causing some baselines to underperform. O2Perm (600 samples) is both heavily right-skewed and long-tailed on a very small dataset, leading most methods to fail on the high-permeability cases.
@@
EA (Electron Affinity, eV) and IP (Ionization Potential, eV). These two datasets conducted from the same SMILES strings with different properties. Data are obtained and processed as Aldeghi and Coley [2]. Figure 4 (e) and (f) show that both properties are well covered and exhibit only moderate skew.
@@ Appendix B Implementation Details
Notice that, we do not add the sparsity penalty until after 50 epochs during the training process. This is to allow the network to freely explore and capture the most salient patterns in the data without prematurely shrinking important weights to zero. This “warm-up” period prevents the model from over-compressing its parameters too early, ensuring it first learns robust representations before enforcing sparsity.
@@
Appendix C Result Details
C.1 More Results on Effectiveness (Q1) and Repetition-Invariant Representation (Q2)
Tables 6 and 7 presents R2 and RMSE on Test5 and Test10 in addition to Test1 and Test60. We can observe that:
@@ C.1 More Results on Effectiveness (Q1) and Repetition-Invariant Representation (Q2)
Algorithm Alignment: Both BFGNN and GRIN-RepAug exhibit a single performance gap between Test1 and Test5, after which their results remain stable through Test10 and Test60. This stability demonstrates the size-generalization benefit conferred by algorithmic alignment. The higher accuracy of GRIN-RepAug compared to BFGNN highlights the additional advantage of alignment with MST.
Algorithm Alignment + Augmentation: GRIN-RepAug’s one-time gap between Test1 and other test sets highlights the crucial role of repeat-unit augmentation in bridging the gap between single RU training and multi-repeat generalization. With augmentation, GRIN maintain near-constant R2 and RMSE across Test1, Test5, Test10, and Test60.
@@
C.2 More Results on Ablation Study
@@
To isolate the effect of the max-aggregator in message passing (see Section 3), we use GRIN-RepAug (i.e., without repeat-unit augmentation) to compare three layer-wise aggregation functions—sum, mean, and max—on GlassTemp and MeltingTemp with 1 repeating unit (Test1) and 60 repeating units (Test60). As shown in Table 9, the max-aggregator consistently delivers the best performance on all test sets. Sum outperforms mean on the Test1 for most cases, while it suffers a much larger performance drop at Test60, underscoring that max-aggregation best preserves size-generalization in the MST-aligned architecture.
@@
To assess the effect of augmentation, we apply RepAug with 3 repeating units to the GCN and GIN baselines and test on GlassTemp and MeltingTemp with 1 repeating unit (Test1) and 60 repeating units (Test60). The results summarized in Tables 10 and 11 show that RepAug yields consistent improvements: on Test1, both GCN+RepAug and GIN+RepAug achieve higher R2 and lower RMSE compared to their non-augmented counterparts. The improvements are more pronounced on Test10, demonstrating that repeat-unit augmentation significantly enhances size-extrapolation performance not only on our GRIN architecture. Figure 2 (a) visualizes the polymer latent representations from GCN and GCN+RepAug with t-SNE.
@@
Appendix D Future Work
While GRIN already establishes a strong foundation for learning repetition-invariant polymer representation, several promising directions could further reinforce its capabilities:
Advanced Algorithm Alignment. Building on our MST-aligned method, future work could explore alignment with more sophisticated dynamic-programming algorithms—such as hierarchical MSTs. This may enable GRIN to learn more comprehensive intra-monomer connections especially for branched or cross-linked polymer architectures.
Broader Material Domains. Our current copolymer dataset comprises two primary architectures—block copolymers, with large contiguous runs of the same monomer, and alternating copolymers, featuring -A-B- repetition at first—each generating distinct structural motifs that affect material properties. Future work could incorporate random copolymers, whose monomer order lacks a fixed pattern, and apply the GRIN framework to each architecture, uncovering architecture-specific invariance and informing tailored representation strategies.

### v2

@@ 4.3 Ablation Study
In this experiment, we evaluate the impact of different training pairs on MeltingTemp (homopolymer) and IP (copolymer). Figure˜3 presents the results. Training pair (with 3-RU augmentation) reaches the best improvement. Further increasing the merge size beyond 3 leads to convergence, with no significant additional gains. These observations are highly compatible with our theory in Section˜3.3.2: the training set lacks binary-update supervision, and the hyperdegree-2 gradient strength plateaus for merge size .
@@
We investigate the effect of different merge ratios in repeat‑unit augmentation on the training set. Since we have demonstrated both theoretically and empirically that a repeat size of three is optimal, this experiment uses the training set . Unsurprisingly, we can observe that in Table˜5, a balanced 1:1 ratio gives the best performance, while reducing the proportion of augmented samples to 1:0.9 and 1:0.8 leads to a monotonic decline in accuracy. This observation further validates our analysis: weakening binary-update supervision (from hyperdegree-2 supernodes) slightly diminishes training effectiveness. More ablation studies are provided in Section˜D.3.
@@ 4.4 Visualization
Figure˜2 visualizes the t-SNE embeddings generated by GCN and GRIN, trained with and without augmentation, evaluated on test sets augmented up to 20 RUs. As discussed in Section˜1, for the same polymer, GCN’s embeddings remain sensitive to repeat size, even with augmentation. In contrast, as shown in Figure˜2 (b), GRIN produces stable embeddings across all test lengths, when combined with 3-RU augmentation, it collapses all repeats of the same polymer into a single cluster. This behavior aligns with the cosine similarity of 0.999 between repeat sizes 1 and 60 reported in Table˜11, demonstrating repetition invariance.
@@ 5 Conclusion
In this work, we addressed limitations of prior polymer learning methods that modeled polymers as single monomers and thus failed to capture their realistic, repetitive structures. We made the first attempt to achieve repetition-invariant representation learning. To this end, we proposed GRIN, which combined MST-aligned aggregation with a repeat-unit-based data augmentation strategy. We provided a theoretical analysis of the minimal augmentation size required to achieve invariance. Experiments on both polymer and copolymer datasets showed that GRIN learned embeddings that reliably generalized to polymers with long repeat lengths.
@@ Appendix A Dataset Details
A.1 Labeled Data
We evaluate our models on four homopolymer property datasets and two copolymer datasets, each split into 60% training, 10% validation, and 30% test sets. To prevent data leakage, we apply repeat‑unit augmentation independently within each split: for every dataset, we generate augmented graphs with repeat sizes from 2 to 60. On the training set, we benchmark four augmentation schemes—{1,2}, {1,3}, {1,2,3}, and {1,4}—while the test set is evaluated on graphs spanning all repeat sizes from 1 through 60.
@@ A.1 Labeled Data
GlassTemp, MeltingTemp, PolyDensity, and O2Perm. Targets are predicting glass transition temperature (GlassTemp, ∘C), polymer density (PolyDensity, g/cm3), melting temperature (MeltingTemp, ∘C) and oxygen permeability (O2Perm, Barrer), respectively. The first three datasets are extracted from Polymer Info [23], O2Perm is compiled from the Membrane Society of Australasia portal following [29].
The distributions displayed in Figure˜4 (a)-(d) indicate that GlassTemp (7000 samples) and MeltingTemp (3600 samples) exhibit only moderate skew, which corresponds to stable performance across all models. PolyDensity (1700 samples) is nearly Gaussian but limited to its smaller size, causing some baselines to underperform. O2Perm (600 samples) is both heavily right‑skewed and long‑tailed on a very small dataset, leading most methods to fail on the high‑permeability cases.
@@
EA (Electron Affinity, eV) and IP (Ionization Potential, eV). These two datasets conducted from the same SMILES strings with different properties. Data are obtained and processed as Aldeghi and Coley [2]. Figure˜4 (e) and (f) show that both properties are well covered and exhibit only moderate skew.
A.2 Unlabeled Data
We leverage 12,764 unlabeled polymers to further evaluate model generalization by generating graph representations, as summarized in Table˜11.
@@ Appendix B Implementation Details
Notice that, we do not add the sparsity penalty until after 50 epochs during the training process. This is to allow the network to freely explore and capture the most salient patterns in the data without prematurely shrinking important weights to zero. This “warm-up” period prevents the model from over‑compressing its parameters too early, ensuring it first learns robust representations before enforcing sparsity.
@@
Appendix C Computational Efficiency
We measured training‐cost breakdown on the MeltingTemp task using a single NVIDIA A6000 GPU. All models were trained under identical hyperparameter settings and employed the same GCN backbone for a fair comparison. Compared to other baselines, GRIN scales efficiently as the training set doubles. GRIN maintains moderate GPU memory usage (557 MiB) and training time (approximately 764 s). It ranks near the middle of the overall efficiency table, faster than multiple baselines such as SSR and only marginally heavier than its lightweight variant, GRIN-RepAug.
| Model | Peak GPU Memory (MiB) | Training Time (s) |
| GCN[15] | 455 | 156.64 |
| IRM[3] | 487 | 250.16 |
| RPGCN[21] | 611 | 424.55 |
| GREA[18] | 587 | 652.27 |
| DIR[35] | 533 | 1419.47 |
| SSR[7] | 531 | 1715.97 |
| DISGEN[14] | 629 | 830.42 |
| BFGCN[22] | 565 | 790.26 |
| GRIN-RepAug | 531 | 438.39 |
| GRIN | 557 | 763.58 |
Appendix D Result Details
D.1 More Results on Effectiveness (Q1)
More Baselines
To evaluate the performance compared to attention- and transformer-based models, which provide stronger long-range interaction modeling and global feature mixing capabilities, we conducted additional experiments using GAT [32] and GraphGPS [25] on the MeltingTemp and O2Perm datasets. As shown in Table˜7, GRIN outperformed both models.
Tables˜8, 9 and 10 presents R2 and RMSE on Test5 and Test10 in addition to Test1 and Test60. We can observe that:
@@ D.1 More Results on Effectiveness (Q1)
Algorithm Alignment: Both BFGNN and GRIN‑RepAug exhibit a single performance gap between Test1 and Test5, after which their results remain stable through Test10 and Test60. This stability demonstrates the size‑generalization benefit conferred by algorithmic alignment. The higher accuracy of GRIN‑RepAug compared to BFGNN highlights the additional advantage of alignment with MST.
Algorithm Alignment + Augmentation: GRIN-RepAug’s one‑time gap between Test1 and other test sets highlights the crucial role of repeat‑unit augmentation in bridging the gap between single RU training and multi‑repeat generalization. With augmentation, GRIN maintain near‑constant R2 and RMSE across Test1, Test5, Test10, and Test60.
@@
Model MeltingTemp O2Perm Test1 Test60 Test1 Test60 R2 RMSE R2 RMSE R2 RMSE R2 RMSE GIN GAT [32] 0.680±0.028 63.9±2.8 0.579±0.029 73.3±2.6 0.873±0.066 760.4±206.9 0.866±0.080 774.1±243.0 GraphGPS [25] 0.650±0.033 66.8±3.2 0.563±0.013 74.7±1.1 0.819±0.168 865.2±408.0 0.754±0.068 1074.1±153.1 GRIN-RepAug 0.741±0.007 57.5±0.8 0.707±0.009 61.1±0.9 0.923±0.008 604.7±31.9 0.910±0.008 655.2±28.6 GRIN 0.745±0.004 57.0±0.4 0.746±0.002 56.9±0.2 0.929±0.002 583.4±7.5 0.929±0.002 581.7±7.2
@@
D.2 More Results on Repetition-Invariant Representation (Q2)
To further evaluate model generalization beyond limited labeled data, we transferred model checkpoints trained on the GlassTemp prediction task to 12,764 unlabeled polymers and quantified repeat-size invariance by computing the cosine similarity between embeddings at repeat sizes 1 and 60. As shown in Table˜11, GRIN practically learns repetition-invariant representation, with a similarity of 0.999 and zero standard deviation.
| Method | Cosine similarity between repeat size 1 and 60 |
| GCN [15] | 0.672±0.119 |
| BFGCN [22] | 0.903±0.060 |
| GRIN | 0.999±0.000 |
D.3 More Results on Ablation Study
@@
To isolate the effect of the max‐aggregator in message passing (see Section˜3), we use GRIN‑RepAug (i.e., without repeat‐unit augmentation) to compare three layer‐wise aggregation functions—sum, mean, and max—on GlassTemp and MeltingTemp with 1 repeating unit (Test1) and 60 repeating units (Test60). As shown in Table˜12, the max‐aggregator consistently delivers the best performance on all test sets. Sum outperforms mean on the Test1 for most cases, while it suffers a much larger performance drop at Test60, underscoring that max‐aggregation best preserves size‑generalization in the MST‐aligned architecture.
@@
To assess the effect of augmentation, we apply RepAug with 3 repeating units to the GCN and GIN baselines and test on GlassTemp and MeltingTemp with 1 repeating unit (Test1) and 60 repeating units (Test60). The results summarized in Tables˜13 and 14 show that RepAug yields consistent improvements: on Test1, both GCN+RepAug and GIN+RepAug achieve higher R2 and lower RMSE compared to their non‑augmented counterparts. The improvements are more pronounced on Test10, demonstrating that repeat‑unit augmentation significantly enhances size‑extrapolation performance not only on our GRIN architecture. Figure˜2 (a) visualizes the polymer latent representations from GCN and GCN+RepAug with t-SNE.
@@
Appendix E Future Work
Limitations
The scale of labeled datasets used in our experiments remains limited, as collecting large, high-quality polymer datasets is inherently challenging. For example, researchers have spent nearly 70 years compiling only about 600 polymers with experimentally measured oxygen permeability in the Polymer Gas Separation Membrane Database [29]. We only partially address the problem by evaluating model generalization on a larger unlabeled polymer dataset by computing representation similarity, as described in appendix˜D. Another limitation is that incorporating augmented training samples in GRIN leads to a modest increase in training time and GPU memory consumption, as reported in appendix˜C.
While GRIN already establishes a strong foundation for learning repetition‑invariant polymer representation, several promising directions could further reinforce its capabilities:
Advanced Algorithm Alignment. Building on our MST-aligned method, future work could explore alignment with more sophisticated dynamic‑programming algorithms, which may enable GRIN to learn more comprehensive intra-monomer connections especially for branched or cross‑linked polymer architectures.
Broader Material Domains. Our current copolymer dataset comprises two primary architectures: block copolymers, with large contiguous runs of the same monomer, and alternating copolymers, featuring -A-B- repetition at first, each generating distinct structural motifs that affect material properties. Future work could incorporate random copolymers, whose monomer order lacks a fixed pattern, and apply the GRIN framework to each architecture, uncovering architecture‑specific invariance and informing tailored representation strategies.